Operations Report – September 5th

MDRS Complete Operations Report 05-SEPT-2021
Name of person filing report: Shannon Rupert
Reason for Report: Beginning of field season/work party daily update
Non-nominal systems: Old leak in the kitchen sink in Hab has resurfaced in the ground below the kitchen
Action taken for non-nominal systems: Nothing to report
Generator: Nothing to report
ScienceDome Dual Split: Operating nominally and I love having AC to escape to when it’s so hot in the afternoon
Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 76.5
Minimum 49
Maximum 100
Note on solar: work party crews are not power conserving, and the heat is adding to our power consumption
Diesel Reading – 0 %.
Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%
Ethanol Free Gasoline – not checked gallons
Water (loft tank): 50 gallons
Water Meter: n/a unit
Water (static tank) – 100 gallons
Static to Loft Pump used – yes
Water in GreenHab – 200 gallons, filled
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 200 gallons
Hab toilet tank emptied: no
Perseverance rover used: Yes
Beginning charge: 100
Ending charge:
Currently charging: yes
Notes on rovers: Nothing to report
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). One of the 350’s is being used as a signReason for use: n/aOil Added? NoATV Fuel Used: Not checked Gals# Hours the ATVs were Used today: 0 Notes on ATVs: Nothing to report
HabCar used and why, where? Yes, hauling water
CrewCar used and why, where? Yes, the team took it to the Goblin Valley this afternoon and it was used at the storage units this morning. It will need to be cleaned before crew use.
Luna used and why, where? No
General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. We will work on radios tomorrow.
Campus-wide inspection, if action taken, what and why?

Volunteer vegetation around campus and MOOP have been removed.
I don’t know if I mentioned yesterday that we outlined in rocks a parking spot for visitors and borrowed the URC orange cones to stop people from coming up to MDRS.
Summary of general operations: We began work on the flooded storage units today but removing anything that has not been used in several years. We announced around town that it was first come, first served on anything outside. We will go tomorrow afternoon and empty more.

The goal is to have everything on shelves, then power wash the mud out, and hopefully have time to move all remaining materials into a single storage unit with nothing on the floor.

Summary of Hab operations: I remembered the leak in the kitchen sink today because the ground is wet again. That will need to be repaired before the first crew arrives.

Summary of Outpost operations: Septic and intern trailer blackwater tanks will be pumped on Thursday.

We finished the construction of the new bed in the intern trailer. It has a double bed on the bottom and a twin on top. They have the best view on Mars.

Summary of GreenHab operations: 60% shade cloth installed. The soil was added to the herb garden, the two greens containers, one of the carrot bins, plus 12 two-gallon pots and 21 one-gallon pots. We made two water runs for the GreenHab today.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Too hot. I am trying to get the team to embrace working in the early morning. So far, they aren’t embracing!

Questions, concerns, supplies needed and requests: Nothing at this time.

Shannon M Rupert, PhD
Mars Desert Research Station
P.O. Box 82
Hanksville, Utah 84734

Copyright © The Mars Society. All rights reserved. | Main Site